# MALADY: Multiclass Active Learning with Auction Dynamics on Graphs

### v2

@@ Abstract
Active learning enhances the performance of machine learning methods, particularly in low-label rate scenarios, by judiciously selecting a limited number of unlabeled data points for labeling, with the goal of improving the performance of an underlying classifier. In this work, we introduce the Multiclass Active Learning with Auction Dynamics on Graphs (MALADY) algorithm which leverages an auction dynamics technique on similarity graphs for efficient active learning. In particular, the proposed algorithm incorporates an active learning loop using as its underlying semi-supervised procedure an efficient and effective similarity graph-based auction method consisting of upper and lower bound auctions that integrate class size constraints. In addition, we introduce a novel active learning acquisition function that incorporates the dual variable of the auction algorithm to measure the uncertainty in the classifier to prioritize queries near the decision boundaries between different classes. Overall, the proposed method can efficiently obtain accurate results using extremely small labeled sets containing just a few elements per class; this is crucial since labeled data is scarce for many applications. Moreover, the proposed technique can incorporate class size information, which improves accuracy even further. Lastly, using experiments on classification tasks and various data sets, we evaluate the performance of our proposed method and show that it exceeds that of comparison algorithms.
@@ 1 Introduction
The choice of training points can have a significant impact on the performance of a machine learning model, particularly in semi-supervised learning (SSL) scenarios where the training set is small. Active learning is a sub-field of machine learning that improves the performance of underlying machine learning methods by carefully selecting unlabeled points to be labeled via the use of a human in the loop or a domain expert. Most active learning methods alternate between (1) training a model using current labeled information and (2) selecting query points from an unlabeled set using an acquisition function that quantifies the utility of each point in the unlabeled set. This iterative process of training a classifier and labeling chosen query points is referred to as the active learning process, whose flowchart is shown in Figure 1. Given a dataset $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N}\}\subset\mathbb{R}^{d}$, we define the labeled set $\mathcal{L}\subset\mathcal{X}$ as the set of labeled points with the corresponding labels $y_{i}=y(x_{i})\in\{1,2,\ldots,C\}$; then, by labeled data, we refer to both the labeled set $\mathcal{L}$ of inputs along with their corresponding labels $\{y_{i}\}_{x_{i}\in\mathcal{L}}$. The unlabeled set then is the set of unlabeled inputs, $\mathcal{U}=\mathcal{X}\setminus\mathcal{L}$. A query set $\mathcal{Q}\subset\mathcal{U}$ is the set of unlabeled points that have been chosen to be labeled and added to the labeled set for the next iteration of the active learning process. While the sets $\mathcal{L}$ and $\mathcal{U}$ change throughout the active learning process as query points are labeled, we simplify our notation of these sets to not explicitly denote the iteration; rather, $\mathcal{L},\ \mathcal{U},$ and $\mathcal{Q}$ will respectively denote the labeled, unlabeled, and query sets at the current iteration of the active learning process.
@@
Given the current $\mathcal{L}$ and $\mathcal{U}$, the main challenge in active learning is to design an acquisition function $\mathcal{A}:\mathcal{U}\rightarrow\mathbb{R}$ that quantifies the benefit of obtaining the label for each currently unlabeled point. The acquisition function values $\mathcal{A}(x)$ for each $x\in\mathcal{U}$ then allow one to prioritize which currently unlabeled data points are chosen to be in the query set $\mathcal{Q}$ to in turn be labeled by a human in the loop or a domain expert. When $|\mathcal{Q}|=1$, the process is referred to as sequential active learning while $|\mathcal{Q}|>1$ corresponds to batch active learning. In this work, we focus on sequential active learning and Figure 1 reflects this specialization. While there are various ways to practically select $\mathcal{Q}\subset\mathcal{U}$ from the set of acquisition function values $\{\mathcal{A}(x)\}_{x\in\mathcal{U}}$, it is most common in sequential active learning to select $\mathcal{Q}=\{x^{\ast}\}$, where $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$ is the maximizer of the acquisition function.
@@
Figure 1: The flowchart of the sequential active learning process. The process starts by training the underlying classifier using the initial labeled set $\mathcal{L}$. A query point $\mathcal{Q}\in\mathcal{U}$ is then selected based on the current acquisition function values. This query point is also labeled according to an oracle (human in the loop or a domain expert) and subsequently, the point is added to the current labeled data $\mathcal{L}$. This process continues until the desired number of points is reached in the labeled set $\mathcal{L}$.
@@
Uncertainty sampling-based acquisition functions [1, 2, 3] are popular for practical active learning and frequently are efficient to compute in application. In particular, uncertainty-based acquisition functions favor query points that are near the current classifier’s decision boundary whose inferred classifications are interpreted as most ”uncertain”. These methods are emblematic of exploitative active learning since they explicitly use measures of distances to decision boundaries to select query points [4]. Moreover, some of the uncertainty-based acquisition functions designed for graph-based classifiers include the uncertainty norm for Poisson reweighted Laplace Learning [4], variance minimization (VOpt) [5], model-change (MC) [6, 7], model-change variance optimality (MCVOpt) [6, 7], $\Sigma$-optimality [8], hierarchical sampling for active learning [9], cautious active clustering [10], and shortest-shortest path $S^{2}$ [11]. In this paper, we introduce a novel uncertainty-based acquisition function that leverages the computed optimal price and incentive values from the underlying auction dynamics graph-based classifier.
@@
Acquisition functions for the active learning process have also been integrated into deep learning frameworks to boost the performance of such parametric models. In particular, some of the recent works include batch active learning by diverse gradient embeddings (BADGE) [12], active learning framework in Bayesian deep learning [13], diverse batch acquisition for deep Bayesian active learning [14], diffusion-based deep active learning [15], stochastic batch acquisition functions [16], and active learning for convolutional neural networks [17].
@@
On the other hand, an important component of the active learning process is the underlying method used for training a classifier based on the currently labeled data $\mathcal{L}$. To this end, (similarity) graph-based methods have proven to be effective models for semi-supervised and active learning settings, particularly in the low label rate regime [18, 19, 20]. These methods leverage both the labeled and unlabeled sets to construct a similarity graph on which the observed labels on $\mathcal{L}$ are used to make inferences of the labels on the unlabeled data according to the graph topology; in this way, the geometric structure of the dataset is modeled with the graph and similar data points receive similar inferred outputs. Such techniques are beneficial since they greatly utilize the structure of and the information from the usually abundant unlabeled data to build a more effective model.
@@
In contrast, variants of auction algorithms, originally developed by Bertsekas [21, 22, 23, 24] for solving the classic assignment problem, have a simple intuitive structure, are easy to code, and have excellent performance in both accuracy and timing. In fact, auction methods outperform substantially their main competitors for important problems, both in theory and in practice, and are also naturally well suited for parallel computation during coding. Therefore, in this paper, we integrate (similarity) graph-based techniques and auction-based procedures to derive the underlying algorithm used in the active learning process, as well as the acquisition function.
@@
In [25], the authors have shown how auction algorithms can be utilized for various applications such as semi-supervised graph-based learning and data classification in the presence of equality or inequality class size constraints on the individual classes. Modifications of this technique for the classification of 3D sensory data are discussed in [26]. Moreover, auction algorithms can be utilized in Merriman, Bence, and Osher (MBO)-based threshold dynamics schemes, which involve threshold dynamics to tackle different tasks; this was shown in [25]. For example, the original MBO scheme [27] was developed in 1992 to simulate and approximate motion by mean curvature; the procedure consists of alternating between diffusion and thresholding. Such an MBO technique was then modified and adapted to a similarity graph setting for the purpose of binary classification and image inpainting in [28], with a multiclass extension developed in [29] and [30]. Applications of the graph-based technique to heat kernel pagerank and hyperspectral imagery and video are detailed in [31] and [32], respectively. Lastly, a summary of recent graph-based optimization approaches for machine learning, including MBO-based techniques, is presented in [33].
@@
With regards to active learning for graph-based semi-supervised classifiers, various recent pipelines have been developed for image processing applications including image segmentation [34], surface water and sediment detection [35], classification of synthetic aperture radar (SAR) data [36, 37], unsupervised clustering of hyperspectal images using nonlinear diffusion [38], hyperspectral unmixing [35], medical image analysis [39], and hyperspectral image classification [40].
@@
  • We propose a novel acquisition function for the active learning process that uses the dual variables of the dual formulations of the modified assignment problems incorporating upper and lower bound class size constraints. The proposed acquisition function not only measures the uncertainty of the volume bound auction algorithm for graph-based semi-supervised problems, but captures salient geometric classifier information that can be leveraged for exploitation of decision boundaries.
@@ 2 Background
In this background section, we first discuss the (similarity) graph construction technique which is fundamental for graph-based methods. Then, we discuss auction dynamics with volume constraints for semi-supervised learning [25] as the technique that is fundamental for the construction of our proposed semi-supervised method. The last subsection reviews some semi-supervised learning algorithms using auction dynamics procedures.
@@ 2.1 Graph Construction
In this section, we review the graph-based framework we use in this paper. Consider a dataset $X=\{x_{1},x_{2},\cdots x_{N}\}\in\mathbb{R}^{d}$, where each data element is represented by a $d$-dimensional feature vector. We generate a undirected graph $G(X,W)$ of vertices and edges between vertices, where $X$ is set of vertices representing the data elements, and $W\in\mathbb{R}^{N\times N}$ is the weight matrix, consisting of weights between pairs of vertices. The weight matrix $W$ is computed using a weight function $w:X\times X\rightarrow\mathbb{R}$, where $W_{i,j}=w(x_{i},x_{j})$ denotes the weight on the edge between the vertices $x_{i}$ and $x_{j}$, where $i\neq j$. Overall, the weight matrix quantifies the similarities between the feature vectors, i.e. data elements, whether labeled or unlabeled; thus, the graph-based framework can provide crucial information about the data and leverages information from the abundant unlabeled data. Some popular weight functions include:
@@
With the assumption that the high dimensional data is concentrated near a low dimensional manifold and the manifold is locally Euclidean, we only compute the $k$ nearest neighbors (kNN) of each of the points using an approximate nearest neighbor search algorithm. This ensures that the weight matrix is sparse and enhances the computational efficiency of our algorithm. For an approximate nearest neighbor search, we use the GraphLearning package [19] which uses the Annoy library [42]. Lastly, to preserve the symmetric property of the weight matrix, we calculate the final weight matrix as ${W^{final}}_{ij}=\{W_{ij}+W_{ji}\}/2$.
@@ 2.2 Auction Algorithm with Exact Volume Constraints
Since our proposed method involves auction techniques originally designed to solve the assignment problem, we review the procedures in this section. Given two disjoint sets $X$ and $Y$ of same cardinality $r$ and a cost function $c:X\times Y\rightarrow\mathbb{R}$, the assignment problem aims to identify a one-to-one correspondence $M=\{(x_{1},y_{1}),\cdots,(x_{r},y_{r})\}$ of the sets $X$ and $Y$, such that the total cost of the matching
@@
$\begin{split}&\max_{\mathbf{z}:X\times Y\to\{0,1\}}\sum_{x\in X}\sum_{y\in Y}c(x,y)z_{y}(x)\\ &\quad\textrm{s.t.}\;\sum_{x\in X}z_{y}(x)=1\;\;\;\forall y,\;\sum_{y\in Y}z_{y}(x)=1\;\;\;\forall x.\end{split}$ (4)
@@
$&\max_{\bm{z\geq 0}}\sum_{x\in X}\sum_{y\in Y}c(x,y)z_{y}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in X}z_{y}(x)=1,\hskip 8.5359pt\sum_{y\in Y}z_{y}(x)=1.\hskip 14.22636pt(LP)\\ \end{aligned}\end{split}$ (5)
This relaxation turns out to be exact, and the problems are equivalent. Furthermore, we can split $Y$ into $N$ similarity classes $\{S_{i}\}_{i=1}^{N}$ each of size $\{V_{i}\}$, and let $c(x,y)=a_{i}(x)$ for each $y\in S_{i}$. With these choices, we can reformulate (5) into:
$&\max_{\bm{z\geq 0}}\sum_{x\in X}\sum_{i=1}^{N}a_{i}(x)\sum_{y\in S_{i}}z_{y}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in X}z_{y}(x)=1,\hskip 8.5359pt\sum_{y\in Y}z_{y}(x)=1\hskip 14.22636pt\\ \end{aligned}\end{split}$ (6)
Using $u_{i}(x)=\sum_{y\in S_{i}}z_{y}(x)$, we obtain the equivalent problem
$&\max_{\bm{u\geq 0}}\sum_{x\in X}\sum_{i=1}^{N}a_{i}(x)u_{i}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in X}u_{i}(x)=S_{i},\hskip 8.5359pt\sum_{i=1}^{N}u_{i}(x)=1.\hskip 14.22636pt\\ \end{aligned}\end{split}$ (7)
We will now focus on solving (7) which is a special case of (5). First, it is helpful to practically interpret (7). In particular, one can view each class in the data classification problem as an institution, such as a gym, that offers a certain number of memberships and the data elements as people trying to obtain a deal on buying access to only one of the institutions. Moreover, the coefficients $a_{i}(x)$ can represent person $x$’s desire to buy the membership in institution $i$. In this view, the solution to (7) maximizes the total satisfaction of the population. Notice that the class size constraints (i.e. number of memberships available) make the assignment of institution memberships to people nontrivial.
One approach to assigning memberships involves the market mechanism. In this case, an institution $j$ is equipped with a price $p_{j}$. Then, person $x$ will want buy the membership from an institution that offers the best value to them:
@@
However, the resulting matching will probably not satisfy the class size constraints. Thus, the challenge is to compute an optimal price vector, called an equilibrium price vector, that results in an institution-person matching which satisfies the constraints on the number of available memberships. The answer lies in the dual formulation of the assignment problem, which can be shown to be the following problem:
$\min_{\bm{p}\in\mathbb{R}^{n},\bm{\pi}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{N}p_{i}S_{i}+\sum_{x}\bm{\pi}(x)$ (9)
@@
In fact, it turns out that the equilibrium price vector $p_{*}$ of the primal problem is in fact the optimal solution of dual problem (9). Moreover, the optimal value of $\bm{\pi}$ is completely determined by $\bm{p}$; given a price vector $\bm{p}$, the above dual problem is minimized when $\pi(x)$ equals the maximum value of $a_{i}(x)-p_{i}$ over $i$.
According to complementary slackness (CS) condition, a complete assignment $M$ and a price vector $\bm{p}$ are primal and dual optimal if and only if each person is assigned to an institution offering the best deal to them. Moreover, the CS condition can be relaxed to allow a person to be assigned to institutions that are within $\epsilon$ of achieving the best deal to them in the definition (8). In particular, we say that an assignment $M$ and a price vector $\bm{p}$ satisfies $\epsilon$-complementary slackness ($\epsilon$-CS) if
@@
In [21], Bertsekas detailed an auction technique for solving the assignment problem by finding the equilibrium (optimal) price vector. Since the original seminal paper, Bertsekas and others have extended the technique to more general problems and have improved the computational aspects of the procedure. For example, [23] develops a technique that efficiently handles assignment problems with multiple identical objects. Overall, an exhaustive reference on auction algorithms is included in [22] which contains information on linear network optimization. Continuous and discrete models for network optimization are described in [24].
Bertsekas’ auction technique contains the following steps: at the beginning of an iteration, one starts with an assignment and price vector satisfying $\epsilon$-CS with $\epsilon>0$. There are two phases in each iteration: the bidding phase and assignment phase; in each iteration, the assignment $M$ and price vector $\bm{p}$ is updated while maintaining $\epsilon$-CS. The two phases are described below:
  • Bidding phase: For each data element $x$, under the assignment $M$:
    • Compute the current value $v_{i}=a_{i}(x)-p_{i}$ for each class $i$, and choose $i^{*}$ such that $i^{*}\in\operatorname*{arg\,max}v_{i}$.
@@
    • Compute the bid $b(x)$ of data element $x$ for class $i^{*}$:

### v3

@@ Abstract
Active learning enhances the performance of machine learning methods, particularly in low-label rate scenarios, by judiciously selecting a limited number of unlabeled data points for labeling, with the goal of improving the performance of an underlying classifier. In this work, we introduce the Multiclass Active Learning with Auction Dynamics on Graphs (MALADY) algorithm, which leverages an auction dynamics technique on similarity graphs for efficient active learning. In particular, the proposed algorithm incorporates an active learning loop using as its underlying semi-supervised procedure an efficient and effective similarity graph-based auction method consisting of upper and lower bound auctions that integrate class size constraints. In addition, we introduce a novel active learning acquisition function that incorporates the dual variable of the auction algorithm to measure the uncertainty in the classifier to prioritize queries near the decision boundaries between different classes. Overall, the proposed method can efficiently obtain accurate results using extremely small labeled sets containing just a few elements per class; this is crucial since labeled data is scarce for many applications. Moreover, the proposed technique can incorporate class size information, which improves accuracy even further. Lastly, using experiments on classification tasks and various data sets, we evaluate the performance of our proposed method and show that it exceeds that of comparison algorithms.
@@ 1 Introduction
The choice of training points can have a significant impact on the performance of a machine learning (ML) model, particularly in semi-supervised learning (SSL) scenarios where the labeled set is small. Active learning (AL) is a sub-field of ML that improves the performance of underlying ML methods by carefully selecting unlabeled points to be labeled via the use of a human in the loop or a domain expert. Most AL methods alternate between (1) training a model using current labeled information and (2) selecting query points from an unlabeled set using an acquisition function that quantifies the utility of each point in the unlabeled set. This iterative process of training a classifier and labeling chosen query points is referred to as the AL process, whose flowchart is shown in Figure 1. Given a dataset $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N}\}\subset\mathbb{R}^{d}$, we define the labeled set $\mathcal{L}\subset\mathcal{X}$ as the set of labeled points with the corresponding labels $y_{i}=y(x_{i})\in\{1,2,\ldots,C\}$; then, by labeled data, we refer to both the labeled set $\mathcal{L}$ of inputs along with their corresponding labels $\{y_{i}\}_{x_{i}\in\mathcal{L}}$. The unlabeled set then is the set of unlabeled inputs, $\mathcal{U}=\mathcal{X}\setminus\mathcal{L}$. A query set $\mathcal{Q}\subset\mathcal{U}$ is the set of unlabeled points that have been chosen to be labeled and added to the labeled set for the next iteration of the AL process. While the sets $\mathcal{L}$ and $\mathcal{U}$ change throughout the active learning process as query points are labeled, we simplify our notation of these sets to not explicitly denote the iteration; rather, $\mathcal{L},\ \mathcal{U},$ and $\mathcal{Q}$ will respectively denote the labeled, unlabeled, and query sets at the current iteration of the AL process.
@@
Given the current $\mathcal{L}$ and $\mathcal{U}$, the main challenge in AL is to design an acquisition function $\mathcal{A}:\mathcal{U}\rightarrow\mathbb{R}$ that quantifies the benefit of obtaining the label for each currently unlabeled point. The acquisition function values $\mathcal{A}(x)$ for each $x\in\mathcal{U}$ then allow one to prioritize which currently unlabeled data points are chosen to be in the query set $\mathcal{Q}$ to in turn be labeled by a human in the loop or a domain expert. When $|\mathcal{Q}|=1$, the process is referred to as sequential AL while $|\mathcal{Q}|>1$ corresponds to batch AL. In this work, we focus on sequential AL, and Figure 1 reflects this specialization. While there are various ways to select $\mathcal{Q}\subset\mathcal{U}$ from the set of acquisition function values $\{\mathcal{A}(x)\}_{x\in\mathcal{U}}$, it is most common in sequential AL to use $\mathcal{Q}=\{x^{\ast}\}$, where $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$ is the maximizer of the acquisition function.
@@
Figure 1: The flowchart of the sequential active learning process. The process starts by training the underlying classifier using the initial labeled set $\mathcal{L}$. A query point $\mathcal{Q}\in\mathcal{U}$ is then selected based on the current acquisition function values. This query point is also labeled according to an oracle (human in the loop or a domain expert), and subsequently, the point is added to the current labeled data $\mathcal{L}$. This process continues until the desired number of points is reached in the labeled set $\mathcal{L}$.
@@
Uncertainty sampling-based acquisition functions [1, 2, 3] are popular for practical active learning and are frequently efficient to compute in applications. In particular, uncertainty-based acquisition functions favor query points that are near the current classifier’s decision boundary whose inferred classifications are interpreted as most ”uncertain”. These methods are emblematic of exploitative AL since they explicitly use measures of distances to decision boundaries to select query points [4]. Moreover, some of the uncertainty-based acquisition functions designed for graph-based classifiers include the uncertainty norm for Poisson-reweighted Laplace Learning [4], variance minimization (VOpt) [5], model-change (MC) [6, 7], model-change variance optimality (MCVOpt) [6, 7], $\Sigma$-optimality [8], hierarchical sampling for active learning [9], cautious active clustering [10], and shortest-shortest path $S^{2}$ [11]. In this paper, we introduce a novel uncertainty-based acquisition function that leverages the computed optimal price and incentive values from the underlying auction dynamics graph-based classifier.
@@
Acquisition functions for the AL process have also been integrated into deep learning frameworks to boost the performance of such parametric models. In particular, some of the recent works include batch AL by diverse gradient embeddings (BADGE) [12], AL framework in Bayesian deep learning [13], diverse batch acquisition for deep Bayesian AL [14], diffusion-based deep AL [15], stochastic batch acquisition functions [16], and AL for convolutional neural networks [17].
@@
On the other hand, an important component of the AL process is the underlying method used for training a classifier based on the currently labeled data $\mathcal{L}$. To this end, (similarity) graph-based methods have proven to be effective models for semi-supervised and AL settings, particularly in the low label rate regime [18, 19, 20, 21]. These methods leverage both the labeled and unlabeled sets to construct a similarity graph on which the observed labels on $\mathcal{L}$ are used to make inferences of the labels on the unlabeled data according to the graph topology; thus, the geometric structure of the dataset is modeled with the graph and similar data points receive similar inferred outputs. Such techniques are beneficial since they utilize the structure of and the information from the usually abundant unlabeled data to build a more effective model.
@@
In contrast, variants of auction algorithms, originally developed by Bertsekas [22, 23, 24, 25] for solving the classic assignment problem, have a simple intuitive structure, are easy to code, and have excellent performance in both accuracy and timing. In fact, auction methods substantially outperform their main competitors for important problems, both in theory and in practice, and are also naturally well-suited for parallel computation during coding. Therefore, in this paper, we integrate (similarity) graph-based techniques and auction-based procedures to derive the underlying algorithm used in the active learning process, as well as the acquisition function.
@@
In [26], the authors have shown how auction algorithms can be utilized for various applications, such as semi-supervised graph-based learning and classification in the presence of equality or inequality class size constraints on the individual classes. Modifications of this technique for the classification of 3D sensory data are discussed in [27]. Moreover, auction methods can be used in Merriman, Bence, and Osher (MBO)-based threshold dynamics schemes, which involve threshold dynamics to tackle different tasks; this was shown in [26]. For example, the original MBO scheme [28] was developed in 1992 to simulate and approximate motion by mean curvature; the procedure consists of alternating between diffusion and thresholding. Such an MBO technique was then modified and adapted to a similarity graph setting for the purpose of binary classification and image inpainting in [29], with a multiclass extension developed in [30] and [31]. Applications of the graph-based technique to heat kernel PageRank and hyperspectral imagery and video are detailed in [32] and [33], respectively. A summary of recent graph-based optimization approaches for machine learning, including MBO-based ones, is presented in [34].
Moreover, graph neural networks (GNN) [35] constitute an important area of modern, graph-based learning methods that have found application in node classification, graph classification, and link prediction. These methods leverage a graph’s structure of nodes and edges to define compositions of transformations of node features in a fashion that is similar to deep neural networks. Notable contributions in this area include [36, 37, 38], with advances in graph classification highlighted in [39, 40].
@@
With regards to active learning for graph-based semi-supervised classifiers, various recent pipelines have been developed for image processing applications, e.g., image segmentation [41], surface water and sediment detection [42], classification of synthetic aperture radar (SAR) data [43, 44], unsupervised clustering of hyperspectral images using nonlinear diffusion [45], hyperspectral unmixing [42], medical image analysis [46], hyperspectral image classification [47], and subgraph matching [48].
@@
  • We propose a novel acquisition function for the active learning process that uses the dual variables of the dual formulations of the modified assignment problems incorporating upper and lower bound class size constraints. The proposed acquisition function not only measures the uncertainty of the volume bound auction algorithm for graph-based semi-supervised problems but also captures salient geometric classifier information that can be leveraged for exploitation of decision boundaries.
@@ 2 Background
In this background section, we first discuss the (similarity) graph construction technique, which is fundamental for graph-based methods. Then, we discuss auction dynamics with volume constraints for semi-supervised learning [26] as the technique that is fundamental for the construction of our proposed semi-supervised method. The last subsection reviews some semi-supervised learning algorithms using auction dynamics procedures. The notation used in this paper is described in Table 1.
Variable Description
$N$ number of nodes
$\mathcal{X}$ dataset, vertex set
$\mathcal{L}$ labeled set
$\mathcal{U}$ unlabeled set
$\mathcal{Q}$ query set
$K$ number of similarity classes
$U_{i}$ upper bound for class $i$
$B_{i}$ lower bound for class $i$
$S_{i}$ similarity class for class $i$
$V_{i}$ cardinality of similarity class $S_{i}$
Table 1: Notation Table
@@ 2.1 Graph Construction
In this section, we review the graph-based framework we use in this paper. Consider a dataset $\mathcal{X}=\{x_{1},x_{2},\cdots x_{N}\}\in\mathbb{R}^{d}$, where each data element is represented by a $d$-dimensional feature vector. We generate a undirected graph $G(\mathcal{X},W)$ of vertices and edges between vertices, where $\mathcal{X}$ is set of vertices representing the data elements, and $W\in\mathbb{R}^{N\times N}$ is the weight matrix, consisting of weights between pairs of vertices. The weight matrix $W$ is computed using a weight function $w:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$, where $W_{i,j}=w(x_{i},x_{j})$ denotes the weight on the edge between the vertices $x_{i}$ and $x_{j}$, where $i\neq j$. Overall, the weight matrix quantifies the similarities between the feature vectors, i.e., data elements, whether labeled or unlabeled; thus, the graph-based framework can provide crucial information about the data and leverage information from the abundant unlabeled data. Some popular weight functions include:
@@
With the assumption that the high-dimensional data is concentrated near a low-dimensional manifold and the manifold is locally Euclidean, we only compute the $k$ nearest neighbors (kNN) of each point using an approximate nearest neighbor search algorithm. This ensures that the weight matrix is sparse and enhances efficiency. For an approximate nearest neighbor search, we use the GraphLearning package [19], which uses the Annoy library [50]. To preserve the symmetric property of the weight matrix, we calculate ${W^{final}}_{ij}=\{W_{ij}+W_{ji}\}/2$.
@@ 2.2 Auction Algorithm with Exact Volume Constraints
Since our proposed method involves auction techniques originally designed to solve the assignment problem, we review the procedures here. Given two disjoint sets $\mathcal{X}$ and $\mathcal{Y}$ of same cardinality $r$ and a benefit function $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the assignment problem aims to identify a one-to-one correspondence $M=\{(x_{1},y_{1}),\cdots,(x_{r},y_{r})\}$ of sets $X$ and $Y$, so that the total benefit of the matching
@@
$\begin{split}&\max_{\mathbf{z}:\mathcal{X}\times\mathcal{Y}\to\{0,1\}}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}c(x,y)z_{y}(x)\\ &\quad\textrm{s.t.}\;\sum_{x\in\mathcal{X}}z_{y}(x)=1\;\;\;\forall y,\;\sum_{y\in\mathcal{Y}}z_{y}(x)=1\;\;\;\forall x.\vspace{-0.5cm}\end{split}$ (4)
@@
$&\max_{\bm{z\geq 0}}\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}c(x,y)z_{y}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in\mathcal{X}}z_{y}(x)=1,\hskip 8.5359pt\sum_{y\in\mathcal{Y}}z_{y}(x)=1.\hskip 14.22636pt(LP)\\ \end{aligned}\end{split}$ (5)
This relaxation turns out to be exact since the solution to a bounded and feasible linear programming problem always includes a vertex of the feasible polytope $\mathcal{P}$. The relaxed constraint set in (5) is the polytope $\mathcal{P}=\left\{\bm{z}\geq 0:\sum_{x\in\mathcal{X}}z_{y}(x)=1,\sum_{y\in\mathcal{Y}}z_{y}(x)=1\right\}$, whose vertices are precisely vectors $\bm{z}$ whose entries are binary. Furthermore, we can split $Y$ into $K$ similarity classes $\{S_{i}\}_{i=1}^{K}$ each of size $V_{i}$, and let $c(x,y)=a_{i}(x)$ for each $y\in S_{i}$. With these choices, we can reformulate (5):
$&\max_{\bm{z\geq 0}}\sum_{x\in\mathcal{X}}\sum_{i=1}^{K}a_{i}(x)\sum_{y\in S_{i}}z_{y}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in\mathcal{X}}z_{y}(x)=1,\hskip 8.5359pt\sum_{y\in\mathcal{Y}}z_{y}(x)=1\hskip 14.22636pt\\ \end{aligned}\end{split}$ (6)
Using $u_{i}(x)=\sum_{y\in S_{i}}z_{y}(x)$, we obtain equivalently:
$&\max_{\bm{u\geq 0}}\sum_{x\in\mathcal{X}}\sum_{i=1}^{K}a_{i}(x)u_{i}(x)\hskip 14.22636pt\\ &\textrm{s.t.}\sum_{x\in\mathcal{X}}u_{i}(x)=V_{i},\hskip 8.5359pt\sum_{i=1}^{K}u_{i}(x)=1.\hskip 14.22636pt\\ \end{aligned}\end{split}$ (7)
We will now focus on solving (7), which is a special case of (5). First, it is helpful to practically interpret (7). In particular, one can view each class in the data classification problem as an institution, such as a gym, that offers a certain number of memberships and the data elements as people trying to obtain a deal on buying access to only one of the institutions. Moreover, the coefficients $a_{i}(x)$ can represent person $x$’s desire to buy the membership in institution $i$. In this view, the solution to (7) maximizes the total satisfaction of the population. Notice that the class size constraints (i.e. number of memberships available) make the assignment of institution memberships to people nontrivial.
One approach to assigning memberships involves the market mechanism where an institution $j$ is equipped with a price $p_{j}$. Then, person $x$ will want to buy the membership from an institution that offers the best value:
@@
However, the resulting matching will probably not satisfy the class size constraints. Thus, the challenge is to compute an optimal price vector, called an equilibrium price vector, that results in an institution-person matching that satisfies the constraints on the number of available memberships. The answer lies in the dual formulation of the assignment problem, which can be shown to be the following problem:
$\min_{\bm{p}\in\mathbb{R}^{K},\bm{\pi}\in\mathbb{R}^{N}}$ $\sum_{i=1}^{K}p_{i}S_{i}+\sum_{x}\bm{\pi}(x)$ (9)
@@
In fact, it turns out that the equilibrium price vector $p_{*}$ of the primal problem is in fact the optimal solution of the dual problem (9). Moreover, the optimal value of $\bm{\pi}$ is completely determined by $\bm{p}$; given a price vector $\bm{p}$, the above dual problem is minimized when $\pi(x)$ equals the maximum value of $a_{i}(x)-p_{i}$ over $i$.
According to the complementary slackness (CS) condition, a complete assignment $M$ and a price vector $\bm{p}$ are primal and dual optimal if and only if each person is assigned to an institution offering the best deal to them. Moreover, the CS condition can be relaxed to allow a person to be assigned to institutions that are within $\epsilon$ of achieving the best deal for them in the definition (8). In particular, we say that an assignment $M$ and a price vector $\bm{p}$ satisfies $\epsilon$-complementary slackness ($\epsilon$-CS) if
@@
In [22], Bertsekas detailed an auction technique for solving the assignment problem by finding the equilibrium (optimal) price vector. Since the seminal paper, Bertsekas and others have extended the technique to more general problems and have improved the computational aspects. For example, [24] develops a technique that efficiently handles assignment problems with multiple identical objects. Overall, an exhaustive reference on auction algorithms is included in [23], which contains information on linear network optimization. Continuous and discrete models for network optimization are described in [25].
Bertsekas’s auction technique contains the following steps: at the beginning of an iteration, one starts with an assignment and price vector satisfying $\epsilon$-CS with $\epsilon>0$. There are two phases in each iteration: the bidding phase and the assignment phase (the assignment $M$ and price vector $\bm{p}$ is updated while maintaining $\epsilon$-CS):
  • Bidding phase: For each $x$, under the assignment $M$:
    • Compute the current value $v_{i}=a_{i}(x)-p_{i}$ for each class $i$, and choose $i^{*}\in\operatorname*{arg\,max}v_{i}$.
@@
    • Compute the bid $b(x)$ of element $x$ for class $i^{*}$: